# Hamiltonian simulation-based quantum-selected configuration interaction for large-scale electronic structure calculations with a quantum computer

### v2

@@ IV Results and Discussion
Subsections A, B, C are devoted to show numerical simulations for the S0 and T1 states of oligoacenes, phenylene-1,4-dinitrene, and hexa-1,2,3,4,5-pentaene, respectively. Subsection D gives a hardware demonstrations for the S0 and T1 states of carbyne molecules.
IV.1 The S0 and T1 states of oligoacenes
@@
The results of the HSB-QSCI simulations of oligoacenes are shown in Figure 4. The RHF and ROHF-like single configurational wave functions are used as the starting wave functions for the Hamiltonian simulations of the S0 and T1 states, respectively, and the number of shots for each time step was set as . The number of Slater determinants in the CAS-CI wave function is 104 (S0) and 61 (T1) for 1, 15912 (S0) and 11076 (T1) for 2, and 2945056 (S0) and 2255121 (T1) for 3, in the point group. Our numerical simulations revealed that the Slater determinants that are important to describe the target electronic states are efficiently sampled from the Hamiltonian simulations. In 1, the system size is very small, and all possible Slater determinants are sampled up to the fifth steps. In 2, chemical precision () was achieved with three steps of the Hamiltonian simulations. In 3, we found that 10 steps of time evolution with shots are not enough to achieve chemical precision. We expect that the number of important Slater determinants increases with , which is the same as the scaling of Hamiltonian terms, when the target state wave function is well approximated by the single determinant as in the ground state of oligoacenes and the energy gap between the ground and the first excited states is unchanged. Considering the fact that the energy gap is smaller in 3 than in 2, we need more than 3.81 times more shots than in the calculation of 2 to obtain the HSB-QSCI energy of 3 with the same accuracy. We have also studied the dependence of the HSB-QSCI energy on the number of shots and the number of time steps in 3. By increasing the number of shots from and simulating more time steps, the number of Slater determinants sampled from the Hamiltonian simulation increases, and the HSB-QSCI energy decreases systematically (see Figure S14 in Supporting Information). Although we need more steps and shots to achieve chemical precision for larger systems, it should be noted that up to 7-electron excited determinants were sampled in a Hamiltonian simulation with in 3. The ability of Hamiltonian simulations to capture higher-order excitations is astounding. These results illustrate the ability of the Hamiltonian simulation to sample a polynomial number of important Slater determinants from exponentially large Hilbert space. When the kernel_fixed_space subroutine in PySCF is used for subspace Hamiltonian diagonalization, chemical precision was achieved with only three steps of time evolutions with shots in 3. However, the HSB-QSCI wave function considers about 57% and 46% of the Slater determinants, hence the HSB-QSCI wave function is less compact (see Figure S11 in Supporting Information).
IV.2 The S0 and T1 states of phenylene-1,4-dinitrene
@@
In the CASSCF optimized orbital basis, the CAS-CI wave function of the S0 state is mainly described by the HF configuration () and the HOMO–LUMO two-electron excited configuration from the HF configuration () (see Table S17 in Supporting Information). In the HSB-QSCI simulations of the S0 state, we performed Hamiltonian simulations with three different initial wave functions, , , and the CAS-CI(2e,2o) wave function , to investigate the dependence of the initial wave function on the energies and the convergence behavior of the HSB-QSCI method. The QSCI simulations were performed under four different conditions: (1) use the results with shots for each time step, (2) use the with shots, (3) use the samples from with shots, and (4) merge the samples from and with shots for each step. For the T1 state, we used the ROHF-like single determinant as the initial wave function for the Hamiltonian simulation. The HSB-QSCI results are summarized in Figure 5. Our numerical simulations suggest that almost the same number of Slater determinants are sampled when and are used. Interestingly, the number of Slater determinants sampled from the Hamiltonian simulation is smaller when is used as the initial wave function. The accuracy of the HSB-QSCI energy with is about an order of magnitude worse than the HF-reference one, but still chemical precision was achieved in 2 steps of time evolution. The singlet–triplet energy gap is calculated to be 1.11 kcal mol-1, in good agreement with the CAS-CI value (1.06 kcal mol-1) and the experimental one [51]. Using PySCF for the subspace Hamiltonian diagonalization slightly increases the number of Slater determinants and yields lower HSB-QSCI energies. The trends of the initial wave function dependence are similar between PyCI- and PySCF-based implementations (see Figure S12 in the Supporting Information).
IV.3 The S0 and T1 states of hexa-1,2,3,4,5-pentaene
@@
The results of the HSB-QSCI simulations are shown in Figure 6. The number of shots for each time step is . In the T1 state calculations, we set the starting wave function to carry two unpaired electrons in the central bonds ( and in Figure 3). It is clear that the number of Slater determinants sampled from the Hamiltonian simulations and the convergence behavior of the HSB-QSCI energies are almost the same for , where is the maximum locality of the qubit Hamiltonian terms. These results are consistent with the trend of the S0 state fidelity value. The singlet–triplet energy gap calculated using the real-time evolution with the untruncated Hamiltonian () is and kcal mol-1 for planar and twisted geometries, respectively, and the difference from the CAS-CI singlet–triplet energy gap is less than 0.1 kcal mol-1 ( = and kcal mol-1 for planar and twisted geometry, respectively). The number of sampled Slater determinants is larger for twisted geometry than for planar geometry, reflecting the spatial distribution of the orbitals, as discussed above. When the maximum locality is set to be = 6 or 8, the error in the HSB-QSCI energy becomes larger. However, our numerical simulations revealed that chemical precision can be achieved even with the maximum locality , by increasing the number of shots to more than and setting the number of maximum time steps to (see Figure S15 in Supporting Information). Note that when using the PySCF kernel_fixed_space subroutine for subspace Hamiltonian diagonalization, chemical precision was achieved with up to five steps of time evolution with shots (see Figure S14 in Supporting Information), but this is a consequence of the quadratic increase of the Slater determinants in PySCF.
@@
It should be emphasized that the ROHF-like single configurational wave function in the LMO basis used in the T1 state calculation has rather a small overlap with the CAS-CI wave function. The squared overlap values in the T1 state are 0.1691 and 0.1240 for planar and twisted geometries, respectively (see Tables S16 and S18 in Supporting Information for the CAS-CI wave function). The fact that the HSB-QSCI provides accurate energy with such a small overlap is promising because the overlap between the HF and the full-CI wave functions decreases with system size, and the preparation of a sophisticated approximate wave function becomes challenging for larger molecules.
IV.4 Hardware demonstration of the HSB-QSCI calculations of the S0 and T1 states of carbyne molecules
@@
Since the Hamiltonian term truncation based on the locality of the qubit Hamiltonian terms works excellently in 5, we performed the HSB-QSCI calculations of 5, 6, and 7 on an IBM Quantum Eagle processor, namely ibm_kawasaki. To reduce the quantum circuit depth of the Hamiltonian simulation, we performed the MPO-based classical compression of the quantum circuit. This approach was demonstrated in our previous study on the quantum phase difference estimation-based energy gap calculations [9]. The details of the MPO-based quantum circuit optimization are given in the reference [9]. In this work, we used a Hamiltonian truncated by the maximum locality ( = 10), and the 10-layer brick wall type quantum circuit is generated for the time evolution operator of . The number of sweeps on the optimizations was set to 10000. In constructing the MPOs, the second-order Trotter decomposition with a single Trotter slice was employed, and the singular value cutoff for the SVD was set to . The order of terms in the Trotter decomposition is set as follows. First, we generate the truncated second quantized Hamiltonian by checking the one- and two-electron integrals in the lexicographic order: after in , and , , , , in that order in . Then the truncated second-quantized Hamiltonian is transformed into a qubit Hamiltonian while retaining the order of the terms, and the Trotterized time evolution operator is constructed subsequently.
@@
| System | CAS-CI | HSB-QSCI; PyCI | HSB-QSCI; PySCF | |||
| #Dets | /Hartree | #Dets | #Dets | |||
| 5 (Planar, S0) | 63504 | 52209 | 0.37 | 63504 | 0.00 | |
| 5 (Planar, T1) | 44100 | 40092 | 0.56 | 44100 | 0.00 | |
| 5 (Twisted, S0) | 63504 | 54585 | 0.33 | 63504 | 0.00 | |
| 5 (Twisted, T1) | 44100 | 40944 | 0.43 | 43260 | 0.01 | |
| 6 (Planar, S0) | 11778624 | 6147941 | 2.00 | 11778624 | 0.00 | |
| 6 (Planar, T1) | 9018009 | 5753210 | 2.93 | 8612604 | 0.01 | |
| 6 (Twisted, S0) | 11778624 | 5838894 | 2.71 | 11758041 | 0.00 | |
| 6 (Twisted, T1) | 9018009 | 5440991 | 2.84 | 8889232 | 0.00 | |
| 7 (Planar, S0) | 2363904400 | 22569820 | 14.26 | 1141628944 | 0.31 | |
| 7 (Planar, T1) | 1914762564 | 14732855 | 23.36 | 526263465 | 3.70 | |
| 7 (Twisted, S0) | 2363904400 | 25866058 | 13.97 | 1096536996 | 0.30 | |
| 7 (Twisted, T1) | 1914762564 | 16356089 | 25.79 | 554484410 | 2.40 | |
The dimensions of the Hamiltonian matrix and the total energies of the HSB-QSCI and the CAS-CI methods are summarized in Table 3. We used and performed shots for each time step in the HSB-QSCI for 6 and 7. In 6, the HSB-QSCI considered about 50–64% of the entire Slater determinant, and it calculated the total energy with an error of about 2.0–2.9 kcal mol-1. Although chemical precision for the total energy was not achieved, the HSB-QSCI provides the singlet–triplet energy gap with an error of 0.93 and 0.12 kcal mol-1 for planar and twisted geometries, respectively, of 6. The energy difference between the planar and twisted geometries is also calculated with an error of about 0.72 and 0.09 kcal mol-1 for the S0 and T1 states, respectively. These results exemplify the ability of the HSB-QSCI to accurately calculate energy differences between two electronic states and two geometries. We expect that the total energy can be further improved by increasing the number of shots in the Hamiltonian simulation. In 7, the error of the HSB-QSCI energy is about 14 kcal mol-1 for the S0 state, and it increases to 23.36 and 25.79 kcal mol-1 for planar and twisted geometries, respectively, of the T1 state. Again, chemical precision for the total energy was not achieved in 7. In addition, the HSB-QSCI overestimated the singlet–triplet energy gap about 9–12 kcal mol-1. As we pointed out above, the CAS-CI wave function of the T1 state in the localized orbital basis shows an inherent multiconfigurational character, and the single Slater determinant used as the initial wave function of the time evolution has small overlap with the CAS-CI wave function. As a result, it is expected that the number of time steps or the number of shots will have to be increased to account for more Slater determinants in order to accurately describe the electronic structure of the T1 state. Nevertheless, it is worth noting that the HSB-QSCI is able to compute the energy difference between the planar and twisted geometries of the S0 state with chemical precision. Note that when PySCF is used for the subspace Hamiltonian diagonalization, the HSB-QSCI provided the total energy of the S0 state with about 0.3 kcal mol-1 of an error, by considering 46–48% of the Slater determinants, but it failed to compute the total energy of the T1 state with chemical precision.
@@
| System | %#Dets222Percentage of the number of Slater determinants calculated as 100 | 333Percentage of the correlation energy in the active space considered in the HSB-QSCI, calculated as 100 |
| 5 (Planar, S0) | 82.21 | 99.99 |
| 5 (Planar, T1) | 90.91 | 99.98 |
| 5 (Twisted, S0) | 85.96 | 99.99 |
| 5 (Twisted, T1) | 92.84 | 99.99 |
| 6 (Planar, S0) | 52.20 | 99.97 |
| 6 (Planar, T1) | 63.80 | 99.96 |
| 6 (Twisted, S0) | 49.57 | 99.96 |
| 6 (Twisted, T1) | 60.33 | 99.96 |
| 7 (Planar, S0) | 0.95 | 99.86 |
| 7 (Planar, T1) | 0.77 | 99.78 |
| 7 (Twisted, S0) | 1.09 | 99.87 |
| 7 (Twisted, T1) | 0.87 | 99.75 |
@@
Table 4 summarizes the percentages of the number of Slater determinants and the correlation energies covered by the HSB-QSCI method in 5–7. It is worth emphasizing that the HSB-QSCI covers more than 99.75% of the correlation energies in the active space in 7, by considering only ca. 1% of all the Slater determinants.

### v3

@@ IV Results and Discussion
Subsections A, B, C, and D are devoted to show numerical simulations for the S0 state of H2O, S0 and T1 states of oligoacenes, phenylene-1,4-dinitrene, and hexa-1,2,3,4,5-pentaene, respectively. Subsection E gives a hardware demonstrations for the S0 and T1 states of carbyne molecules.
IV.1 The S0 state of the H2O molecule
In this subsection, we discuss the performance of the sampling strategies for Slater determinants in the QSCI, using the H2O molecule. We compare the HSB-QSCI, QSCI with randomly sampled Slater determinants that preserve spatial symmetry, , and , VQE-based QSCI [14], and HCI. HSB-QSCI calculations were performed with , and the number of shots per time step was set to . For the random sampling-based QSCI, we always included the HF configuration in the subspace Hamiltonian, and excited determinants belonging to the same irreducible representation of the S0 state and having the correct and . The number of excited determinants satisfying these conditions is 99. We performed five independent simulations. The HCI is performed using the PyCI package [55]. The results are summarized in Figure 4. In HSB-QSCI with , the energy error is already less than Hartree (0.36 kcal mol-1). Chemical precision is achieved with only one time evolution step, considering only 20 Slater determinants. This is in contrast to the VQE-based QSCI [14], which requires about 50 VQE iterations and 32 Slater determinants to achieve an energy error below Hartree. Note that they also reported that 16 Slater determinants are enough to achieve chemical precision, after VQE convergence. It is worth noting that the probability of sampling the HF configuration in the first time step of HSB-QSCI is approximately 0.942. In contrast, in the perfect state preparation limit of VQE-based sampling, this probability is calculated to be 0.978. This indicates that HSB-QSCI has a higher chance of sampling excited determinants than the VQE-based approach. Of course, the quantum circuit is deeper in HSB-QSCI than in the VQE-based one. However, improvements in the overall number of required shots and the compactness of the QSCI wave function are significant. HCI yields slightly lower energy than HSB-QSCI with 20 Slater determinants, Hartree = 0.032 kcal mol-1. However, as the number of Slater determinants increases, the difference in energy errors between HSB-QSCI and HCI diminishes, and the two methods give identical energies when 28 Slater determinants are used. It is important to note that HSB-QSCI yields energies comparable to HCI, although starting from an approximate wave function in HSB-QSCI may lead to sampling Slater determinants that are more relevant to excited states than to the ground state. The comparison with the random sampling-based implementation is even more illuminating. In this case, the energy error remains large even when 80% of the Slater determinants are included in the subspace Hamiltonian diagonalization.
IV.2 The S0 and T1 states of oligoacenes
@@
The results of the HSB-QSCI simulations of oligoacenes are shown in Figure 5. The RHF and ROHF-like single configurational wave functions are used as the starting wave functions for the Hamiltonian simulations of the S0 and T1 states, respectively, and the number of shots for each time step was set as . The number of Slater determinants in the CAS-CI wave function is 104 (S0) and 61 (T1) for 1, 15912 (S0) and 11076 (T1) for 2, and 2945056 (S0) and 2255121 (T1) for 3, in the point group. Our numerical simulations revealed that the Slater determinants that are important to describe the target electronic states are efficiently sampled from the Hamiltonian simulations. In 1, the system size is very small, and all possible Slater determinants are sampled up to the fifth steps. In 2, chemical precision () was achieved with three steps of the Hamiltonian simulations. In 3, we found that 10 steps of time evolution with shots are not enough to achieve chemical precision. We expect that the number of important Slater determinants increases with , which is the same as the scaling of Hamiltonian terms, when the target state wave function is well approximated by the single determinant as in the ground state of oligoacenes and the energy gap between the ground and the first excited states is unchanged. Considering the fact that the energy gap is smaller in 3 than in 2, we need more than 3.81 times more shots than in the calculation of 2 to obtain the HSB-QSCI energy of 3 with the same accuracy. We have also studied the dependence of the HSB-QSCI energy on the number of shots and the number of time steps in 3. By increasing the number of shots from and simulating more time steps, the number of Slater determinants sampled from the Hamiltonian simulation increases, and the HSB-QSCI energy decreases systematically (see Figure S14 in Supporting Information). Although we need more steps and shots to achieve chemical precision for larger systems, it should be noted that up to 7-electron excited determinants were sampled in a Hamiltonian simulation with in 3. It is known that higher-order excitations are important to compute the energy in a quantitative manner, especially in strongly correlated systems [19]. The ability of Hamiltonian simulations to capture higher-order excitations is remarkable and is an important feature for applications to large and strongly correlated systems. These results illustrate the ability of the Hamiltonian simulation to sample a polynomial number of important Slater determinants from exponentially large Hilbert space. Note that similar trend has also been reported by other groups [60, 61]. When the kernel_fixed_space subroutine in PySCF is used for subspace Hamiltonian diagonalization, chemical precision was achieved with only three steps of time evolutions with shots in 3. However, the HSB-QSCI wave function considers about 57% and 46% of the Slater determinants, hence the HSB-QSCI wave function is less compact (see Figure S11 in Supporting Information).
IV.3 The S0 and T1 states of phenylene-1,4-dinitrene
@@
In the CASSCF optimized orbital basis, the CAS-CI wave function of the S0 state is mainly described by the HF configuration () and the HOMO–LUMO two-electron excited configuration from the HF configuration () (see Table S17 in Supporting Information). In the HSB-QSCI simulations of the S0 state, we performed Hamiltonian simulations with three different initial wave functions, , , and the CAS-CI(2e,2o) wave function , to investigate the dependence of the initial wave function on the energies and the convergence behavior of the HSB-QSCI method. The QSCI simulations were performed under four different conditions: (1) use the results with shots for each time step, (2) use the with shots, (3) use the samples from with shots, and (4) merge the samples from and with shots for each step. For the T1 state, we used the ROHF-like single determinant as the initial wave function for the Hamiltonian simulation. The HSB-QSCI results are summarized in Figure 6. Our numerical simulations suggest that almost the same number of Slater determinants are sampled when and are used. Interestingly, the number of Slater determinants sampled from the Hamiltonian simulation is smaller when is used as the initial wave function. The accuracy of the HSB-QSCI energy with is about an order of magnitude worse than the HF-reference one, but still chemical precision was achieved in 2 steps of time evolution. Thus, using the CAS-CI wave function with a smaller active space can help generate a more compact HSB-QSCI wave function for strongly correlated systems. However, preparing the CAS-CI reference state itself may become a bottleneck when the number of quasi-degenerate orbitals is large. One possible solution to this challenge is to adopt a technique for constructing a multiconfigurational wave function without performing post-HF calculations [67]. The singlet–triplet energy gap is calculated to be 1.11 kcal mol-1, in good agreement with the CAS-CI value (1.06 kcal mol-1) and the experimental one [65]. Using PySCF for the subspace Hamiltonian diagonalization slightly increases the number of Slater determinants and yields lower HSB-QSCI energies. The trends of the initial wave function dependence are similar between PyCI- and PySCF-based implementations (see Figure S12 in the Supporting Information).
IV.4 The S0 and T1 states of hexa-1,2,3,4,5-pentaene
@@
The results of the HSB-QSCI simulations are shown in Figure 7. The number of shots for each time step is . In the T1 state calculations, we set the starting wave function to carry two unpaired electrons in the central bonds ( and in Figure 3). It is clear that the number of Slater determinants sampled from the Hamiltonian simulations and the convergence behavior of the HSB-QSCI energies are almost the same for , where is the maximum locality of the qubit Hamiltonian terms. These results are consistent with the trend of the S0 state fidelity value. The singlet–triplet energy gap calculated using the real-time evolution with the untruncated Hamiltonian () is and kcal mol-1 for planar and twisted geometries, respectively, and the difference from the CAS-CI singlet–triplet energy gap is less than 0.1 kcal mol-1 ( = and kcal mol-1 for planar and twisted geometry, respectively). The number of sampled Slater determinants is larger for twisted geometry than for planar geometry, reflecting the spatial distribution of the orbitals, as discussed above. When the maximum locality is set to be = 6 or 8, the error in the HSB-QSCI energy becomes larger. However, our numerical simulations revealed that chemical precision can be achieved even with the maximum locality , by increasing the number of shots to more than and setting the number of maximum time steps to (see Figure S15 in Supporting Information). In this case, we observed an abrupt improvement in energy convergence. Since we did not observe such behavior when the untruncated Hamiltonian was used for the time evolution, we suspect that Hamiltonian truncation is responsible for this. When a Hamiltonian truncated to maximum locality is used for time evolution, the probability of measuring excited determinants with long-range excitations becomes smaller, because long-range interaction terms are discarded by the truncation. Consequently, the use of the truncated Hamiltonian in the time evolution operator causes an imbalance in the description of short-range and long-range interactions, which is likely responsible for the observed behavior.
Note that when using the PySCF kernel_fixed_space subroutine for subspace Hamiltonian diagonalization, chemical precision was achieved with up to five steps of time evolution with shots (see Figure S14 in Supporting Information), but this is a consequence of the quadratic increase of the Slater determinants in PySCF.
@@
It should be emphasized that the ROHF-like single configurational wave function in the LMO basis used in the T1 state calculation has a rather small overlap with the CAS-CI wave function. The squared overlap values in the T1 state are 0.1691 and 0.1240 for planar and twisted geometries, respectively (see Tables S16 and S18 in Supporting Information for the CAS-CI wave function). The fact that the HSB-QSCI provides accurate energy with such a small overlap is promising because the overlap between the HF and the full-CI wave functions decreases with system size, and the preparation of a sophisticated approximate wave function becomes challenging for larger molecules.
IV.5 Hardware demonstration of the HSB-QSCI calculations of the S0 and T1 states of carbyne molecules
@@
Since the Hamiltonian term truncation based on the locality of the qubit Hamiltonian terms works excellently in 5, we performed the HSB-QSCI calculations of 5, 6, and 7 on an IBM Quantum Eagle processor, namely ibm_kawasaki. To reduce the quantum circuit depth of the Hamiltonian simulation, we performed the MPO-based classical compression of the quantum circuit. This approach was demonstrated in our previous study on the quantum phase difference estimation-based energy gap calculations [9]. The details of the MPO-based quantum circuit optimization are given in the reference [9]. In this work, we used a Hamiltonian truncated by the maximum locality ( = 10), and the 10-layer brick wall type quantum circuit is generated for the time evolution operator of in atomic unit. The number of sweeps on the optimizations was set to 10000. In constructing the MPOs, the second-order Trotter decomposition with a single Trotter slice was employed, and the singular value cutoff for the SVD was set to . The order of terms in the Trotter decomposition is set as follows. First, we generate the truncated second quantized Hamiltonian by checking the one- and two-electron integrals in the lexicographic order: after in , and , , , , in that order in . Then the truncated second-quantized Hamiltonian is transformed into a qubit Hamiltonian while retaining the order of the terms, and the Trotterized time evolution operator is constructed subsequently.
@@
| System | HF | CAS-CI | HSB-QSCI | |||
| /Hartree | #Dets | /Hartree | %#Detsa | %b | ||
| 5 (Planar, S0) | 63504 | 82.21 | 0.37 | 99.98 | ||
| 5 (Planar, T1) | 44100 | 90.91 | 0.56 | 99.97 | ||
| 5 (Twisted, S0) | 63504 | 85.96 | 0.33 | 99.98 | ||
| 5 (Twisted, T1) | 44100 | 92.84 | 0.43 | 99.99 | ||
| 6 (Planar, S0) | 11778624 | 52.20 | 2.00 | 99.92 | ||
| 6 (Planar, T1) | 9018009 | 63.80 | 2.93 | 99.88 | ||
| 6 (Twisted, S0) | 11778624 | 49.57 | 2.71 | 99.89 | ||
| 6 (Twisted, T1) | 9018009 | 60.33 | 2.84 | 99.89 | ||
| 7 (Planar, S0) | 2363904400 | 0.95 | 14.26 | 99.55 | ||
| 7 (Planar, T1) | 1914762564 | 0.77 | 23.36 | 99.27 | ||
| 7 (Twisted, S0) | 2363904400 | 1.09 | 13.97 | 99.56 | ||
| 7 (Twisted, T1) | 1914762564 | 0.87 | 25.79 | 99.18 | ||
a Percentage of the number of Slater determinants calculated as 100
@@
b Percentage of the correlation energy in the active space considered in the HSB-QSCI, calculated as 100 .
The dimensions of the Hamiltonian matrix and the total energies of the HSB-QSCI and the CAS-CI methods are summarized in Table 3. We used and performed shots for each time step in the HSB-QSCI for 6 and 7. In 6, the HSB-QSCI considered about 50–64% of the entire Slater determinant, and it calculated the total energy with an error of about 2.0–2.9 kcal mol-1. Although chemical precision for the total energy was not achieved, the HSB-QSCI provides the singlet–triplet energy gap with an error of 0.93 and 0.12 kcal mol-1 for planar and twisted geometries, respectively, of 6. The energy difference between the planar and twisted geometries is also calculated with an error of about 0.72 and 0.09 kcal mol-1 for the S0 and T1 states, respectively. These results exemplify the ability of the HSB-QSCI to accurately calculate energy differences between two electronic states and two geometries. We expect that the total energy can be further improved by increasing the number of shots in the Hamiltonian simulation. In 7, the error of the HSB-QSCI energy is about 14 kcal mol-1 for the S0 state, and it increases to 23.36 and 25.79 kcal mol-1 for planar and twisted geometries, respectively, of the T1 state. Again, chemical precision for the total energy was not achieved in 7. In addition, the HSB-QSCI overestimated the singlet–triplet energy gap about 9–12 kcal mol-1. As we pointed out above, the CAS-CI wave function of the T1 state in the localized orbital basis shows an inherent multiconfigurational character, and the single Slater determinant used as the initial wave function of the time evolution has small overlap with the CAS-CI wave function. As a result, it is expected that the number of time steps or the number of shots will have to be increased to account for more Slater determinants in order to accurately describe the electronic structure of the T1 state. Nevertheless, it is worth noting that the HSB-QSCI is able to compute the energy difference between the planar and twisted geometries of the S0 state with chemical precision. Note that when PySCF is used for the subspace Hamiltonian diagonalization, the HSB-QSCI provided the total energy of the S0 state with about 0.3 kcal mol-1 of an error, by considering 46–48% of the Slater determinants, but it failed to compute the total energy of the T1 state with chemical precision. See Table S23 in Supporting Information for details.
@@
Table 3 also summarizes the percentages of the number of Slater determinants and the correlation energies covered by the HSB-QSCI method in 5–7. Here, we calculated the correlation energies as the energy difference between HF (RHF and ROHF for the S0 and T1 state, respectively, in the canonical orbital basis) and the HSB-QSCI energies. It is worth emphasizing that the HSB-QSCI covers more than 99.18% of the correlation energies in the active space in 7, by considering only ca. 1% of all the Slater determinants.